The circle whose center is located in the first coordinate quarter touches the axis Ox at the point M, intersects two hyperbolas y=k1/x and y=k2/x (k1,k2 > 0) at the points A and B such that the line AB passes through the origin O. It is known that (4/k1) + (1/k2) = 20. Find the smallest possible length of the OM segment. In response, write down the square of the length of the segment OM.
1
Expert's answer
2020-12-20T18:04:30-0500

"|OM|^2=|OA|\cdot|OB|"

Points "A(x_A,y_A)" and "B(x_B,y_B)" are points of intersection circle with hyperbolas.

"|OM|^2=\sqrt{(x_A^2+y_A^2)(x_B^2+y_B^2)}"

"|OM|^2=\sqrt{(x_A^2+(k_1/x_A)^2)(x_B^2+(k_2/x_B)^2)}"

Since A and B lie on the same line, then: "x_A\/y_A=x_B\/y_B"

"\frac{x_Ay_A}{x_By_B}=\frac{k_1}{k_2}\implies (\frac{x_A}{x_B})^2=\frac{k_1}{k_2}"


"|OM|^2=\sqrt{\frac{(x_A^4+k_1^2)(x_B^4+k_2^2)}{x_A^2x_B^2}}=\sqrt{x_A^2x_B^2+2k_1k_2+k_1^2k_2^2/x_A^2x_B^2}"

"x_B^2=\frac{k_2}{k_1}x_A^2"

"|OM|^2==\sqrt{\frac{k_2}{k_1}x_A^4+2k_1k_2+k_1^3k_2/x_A^4}"

"(4\/k1) + (1\/k2) = 20"

"k_2=\frac{k_1}{20k_1-4}"

"|OM|^2==\sqrt{\frac{k_1x_A^8+2x_A^4k_1^3+k_1^5}{x_A^4k_1(20k_1-4)}}"


"|OM|^2=\frac{1}{x_A^2}\sqrt{\frac{x_A^8+2x_A^4k_1^2+k_1^4}{20k_1-4}}"


Let: "x=x_A, k=k_1, |OM|=y(x)"

Then:

"y(x)=\frac{1}{x}(\frac{x^8+2x^4k^2+k^4}{20k-4})^{1/4}"

"y'(x)=\frac{1}{4}(\frac{x^4+2k^2}{20k-4}+\frac{k^4}{x^4(20k-4)})^{-3/4}\cdot(\frac{4x^3}{20k-4}-\frac{4k^4}{x^5(20k-4)})=0"

"4x^8-4k^4=0"

"x=\sqrt{k}"

"x_A=\sqrt{k_1}"


"|OM|_{min}^2=\frac{1}{k_1}\sqrt{\frac{k_1^4+2k_1^4+k_1^4}{20k_1-4}}=\frac{k_1}{\sqrt{5k_1-1}}"